OpenStudy (blackstreet23):

Divide 36500 by 0.00900 and enter the answer using the proper scientific "E" notation so that the result is within 1% of the exact calculation.

OpenStudy (blackstreet23):

How do to this?

OpenStudy (blackstreet23):

@Shalante

OpenStudy (anonymous):

What did you get when you divide it? 2.7e^3=2.7 x10^3 =2700

OpenStudy (blackstreet23):

4055555.55556

OpenStudy (blackstreet23):

I got this

OpenStudy (blackstreet23):

would it be 4.05e+11 ?

OpenStudy (blackstreet23):

@Shalante

OpenStudy (blackstreet23):

https://www.google.com/?gws_rd=ssl#q=36500+%2F0.00900

OpenStudy (blackstreet23):

i just divided it by google

OpenStudy (anonymous):

Yes, that is correct I was eating dinner, my bad.

OpenStudy (blackstreet23):

@freckles

OpenStudy (blackstreet23):

@SolomonZelman

OpenStudy (blackstreet23):

@amistre64

OpenStudy (amistre64):

36500 ------ 0.00900 365(100000) ----------- 9 40.555... (10^5) 4055555. 555... what does it mean to be within 1% of the actual value?

OpenStudy (amistre64):

by the way .55555... = 5/9

OpenStudy (blackstreet23):

got it right ! it was : (4055555.55556*0.01)+ 4055555.55556 (4055555.55556*0.01- 4055555.55556 4015000 <4055555.55556<4096111.11112

OpenStudy (blackstreet23):

Thanks for taking the time and helping me :D

OpenStudy (amistre64):

:) good job the

OpenStudy (amistre64):

**then

OpenStudy (blackstreet23):

4.06 e+6 was the right answer